# Hypocoercivity and metastability of degenerate KFP equations at low temperature

### v1

@@ Abstract.
We consider Kramers-Fokker-Planck operators with general degenerate coefficients. We prove semiclassical hypocoercivity estimates for a large class of such operators. Then, we manage to prove Eyring–Kramers formulas for the bottom of the spectrum of some particular degenerate operators in the semiclassical regime, and quantify the spectral gap separating these eigenvalues from the rest of the spectrum. The main ingredient is the construction of sharp Gaussian quasimodes through an adaptation of the WKB method.
@@ 1.1. Motivations
with drift $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and diffusion matrix $\sigma:\mathbb{R}^{d}\to{\mathcal{M}}_{d}(\mathbb{R})$, it is defined as
@@
This operator is called the (semiclassical) Kramers-Fokker-Planck operator and takes the form
@@
One of the pioneer work studying $P_{KFP}$ is [19] in which the authors determined estimates regarding the return to equilibrium for the semigroup $e^{-tP_{KFP}}$ for certain potentials that are homogeneous near infinity. This result has then been generalized for example in [27, 2]. In [20] the authors derived a rough localization of the small eigenvalues of $P_{KFP}$ in the semiclassical limit $h\to 0$. More precise asymptotics have been obtained in [17, 18]. Note that these work are not restricted to the KFP operator, the authors considered a wider class of supersymmetric non-self-adjoint operators. Then, this supersymmetric assumption has been relaxed in [5]. The goal of this paper is to obtain similar result for a KFP operator with general coefficients that does not fit the assumptions of [17, 18, 5].
@@ 1.2. Main results
However, the kinetic structure of $P$ is very important and yields powerful tools to palliate [17]. We can make use of the notion of hypocoercivity. This notion, initiated in [21], and developed further in [33, 7] and references therein. Here we are in a setting which is very convenient to use a semiclassical version of [9] and we shall follow this strategy to obtain the desired resolvent estimates.
Throughout the paper, we shall consider a potential satisfying
@@
We denote ${\mathcal{U}}$ the set of critical points of $V$, and we consider the partition ${\mathcal{U}}={\mathcal{U}}^{odd}\sqcup{\mathcal{U}}^{even}$ where $x^{*}\in{\mathcal{U}}^{even}\iff\forall i,\ \nu_{i}^{x^{*}}\in 2\mathbb{N}$. Now we shall say that $x^{*}\in{\mathcal{U}}^{even}$ is of index $j\in\operatorname{[\negthinspace[}0,d\operatorname{]\negthinspace]}$ if $\sharp\{i\ |\ t_{i}^{x^{*}}<0\}=j$ and therefore ${\mathcal{U}}^{even}=\bigsqcup_{j=0}^{d}{\mathcal{U}}^{(j)}$ where ${\mathcal{U}}^{(j)}$ is the set of critical points of $V$ with even order in each direction and of index $j$, we also denote $n_{0}=\sharp{\mathcal{U}}^{(0)}$ the number of minima of $V$.
Notice that the critical points of $f$ are exactly the $(x^{*},0)$, with $x^{*}$ a critical point of $V$, Moreover we have the same partition ${\mathcal{U}}={\mathcal{U}}^{odd}\sqcup{\mathcal{U}}^{even}$ and the critical points have the same index. In the following we will identify those two and use $x^{*}$ instead of $(x^{*},0)$ where it is clear which one we are really talking about ($x^{*}$ will mostly be denoted either $\mathbf{m}$ if of index $0$ or $\mathbf{s}$ if of index $1$).
In the following we denote
(1.15) $\overline{\nu}=\max_{i,x^{*}}\nu_{i}^{x^{*}}$
@@
There exists $g_{1},g_{2}>0$ such that for all $x\in\mathbb{R}^{d}$,
@@
Suppose Assumptions 1, 2, 5, 6 and 9 hold true. There exists $h_{0}>0$, $c_{0},c_{1},c>0$, such that for all $h\in]0,h_{0}]$, there exists $G_{h}$ subspace of $L^{2}(\mathbb{R}^{d+d^{\prime}})$ of dimension $n_{0}$ such that
@@
then there exists $\lambda_{\mathbf{m}}(h)\in\mathbb{C}$ for all $\mathbf{m}\in{\mathcal{U}}^{(0)}$ such that $\sigma(P)\cap\{|\operatorname{Re}z|\leq c_{0}g(h)\}=\{\lambda_{\mathbf{m}}(h),\mathbf{m}\in{\mathcal{U}}^{(0)}\}$ counted with multiplicity, and for all $\mathbf{m}\in{\mathcal{U}}^{(0)},\ |\lambda_{\mathbf{m}}(h)|\leq e^{-c/h}$. And (still under (1.17)), for all $0<c_{0}^{\prime}<c_{1}$
@@
For the rough localization of Theorem 1, we only need vague constructions around the critical points of $V$ in order to have Proposition 2.1. In order to obtain sharp asymptotics, we have to refine our constructions. It requires the introduction of the topological definitions we recall in Section A.2.
@@
The function $g$ defined in (2.18) is no worse than polynomial. That is, there exists $h_{0},c>0$ such that for all $h\in(0,h_{0}]$,
@@
Suppose the hypothesis of Theorem 1, Assumptions 7, 8, and (Gener) hold true. Suppose moreover that Proposition 4.1 apply. There exists $h_{0},\gamma>0$ such that for all $h\in\ ]0,h_{0}]$, one has $\lambda(\underline{\mathbf{m}},h)=0$ and for all $\mathbf{m}\neq\underline{\mathbf{m}}$, $\lambda(\mathbf{m},h)$ satisfies the following Eyring–Kramers type formula
@@
As mentioned in the paragraph after (Gener), it seems that this assumption is not required in order to obtain sharp results although we did no computation if it does not hold. For Assumption 8, we observe that it is very commonly satisfied and all examples in this article satisfy it.
@@ 1.3. Metastability
Furthermore, we can describe the metastable behavior of solutions of (1.18).
@@
In the setting of Theorem 2, let $S_{1}\leq\cdots\leq S_{p+1}=+\infty$ denote the non-decreasing sequence of the $S(\mathbf{m})$’s defined in (A.1) such that if $S_{k}=S_{k+1}$, then $\mu_{k}\leq\mu_{k+1}$ and let $\Pi^{\leq}_{k}$ be the spectral projector of $P$ associated with its eigenvalues of modulus less than $h^{\mu_{k}}e^{-2S_{k}/h}$. For two positive functions $t_{\pm}(h)$ such that $t_{-}(h)=O(h^{\infty})$ and $t_{+}^{-1}(h)=o(h|\ln h|^{-1})$, we define the times
$t_{0}^{+}=t_{+}(h)\ \ \text{ and }\forall 1\leq k\leq p+1,\ t^{\pm}_{k}=t_{\pm}(h)e^{2S_{k}/h}$
@@
We can link this with the stochastic process solving (1.5). Under mild hypotheses, one can show that the probability density $\rho(t,\cdot)$ of the process $(X_{t})$ solution of (1.5) is solution to the problem
@@ Acknowledgements
Then, we want to obtain precise Eyring-Kramers law for our operator. We did not manage to obtain such result in broad generality, but we could prove these in several prescribed situations. Section 4 focuses on the local constructions of the WKB method in the spirit of [5], while the purpose of Section 5 is to glue these local constructions to obtain globally defined cutoffs. This leads to the proof of Theorem 2.
@@ 2. Hypocoercive estimates
The following assumption will help us prove some bounds on $A$ for the hypocoercivity result
@@
    $\forall(x,v)\in\mathbb{R}^{d+d^{\prime}}\ \int\alpha(x,\sqrt{h}v)e^{-2|\Sigma v|^{2}}\ dv=0$
@@
    where we denote $J_{y}\gamma$ the Jacobian matrix of $\gamma$ with respect to the variable $y$ and $\Delta_{v}\alpha$ the vector $(\Delta_{v}\alpha_{i})_{i}$.
@@ 2.1. Proof of Theorem 1
With the expression of $\delta(h)$ and Lemma 2.4, let us denote for $h>0$
@@ 5.1. Proof of Theorem 2 and graded matrices
$e^{-tP/h}=\Pi_{0}+O(e^{-t\underset{\mathbf{m}\neq\underline{\mathbf{m}}}{\min}\operatorname{Re}(\lambda(\mathbf{m},h))(1-h^{N})/h}),$
@@
which is a bounded operator, we have $\Pi_{0}=e^{-tP/h}\Pi_{0}$ and it suffices to show that for all $N\in\mathbb{N}$,
@@
and $D_{p+1}=D(0,R^{-1}h^{\mu_{p}}e^{-2S_{p}/h})$. For $R$ fixed large enough and every $h$ small enough, each exponentially small eigenvalue of $P$ belongs to exactly one of the disjoint sets $D_{k}$ from Theorem 2. Moreover, $\partial D_{k}$ is at distance of order $h^{\mu_{k}}e^{-2S_{k}/h}$ (resp. $h^{\mu_{p}}e^{-2S_{p}/h}$) from the spectrum of $P$ for $k\in\operatorname{[\negthinspace[}1,p\operatorname{]\negthinspace]}$ (resp. $k=p+1$). Using the resolvent estimate of $P$ on the image of $\Pi_{\mathscr{C}}$ given by Theorem 3 and (5.18) to control the contribution of on the image of $1-\Pi_{\mathscr{C}}$, we get
@@
using that $\mu_{k}\geq 1$. On the contrary, for $t_{k}^{+}\leq t$, (5.21) and $e^{-t_{k}^{+}h^{\mu_{k}-1}e^{-2S_{k}/h}/R}=O(h^{\infty})$ (using that $\mu_{k}\leq 2$ from its definition) give
@@
Lastly, $e^{-tP/h}\Pi_{p+1}=\Pi_{p+1}$ since $\Pi_{p+1}$ is the rank-one spectral projector associated with the eigenvalue $0$. On the other hand, since $e^{-\varepsilon t_{0}^{+}}=O(h^{\infty})$, (5.20) becomes

### v2

@@ Abstract.
We consider Kramers-Fokker-Planck operators with general degenerate coefficients. We prove semiclassical hypocoercivity estimates for a large class of such operators. Then, we manage to prove Eyring-Kramers formulas for the bottom of the spectrum of some particular degenerate operators in the semiclassical regime, and quantify the spectral gap separating these eigenvalues from the rest of the spectrum. The main ingredient is the construction of sharp Gaussian quasimodes through an adaptation of the WKB method.
@@ 1.1. Motivations
with drift $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and diffusion matrix $\sigma:\mathbb{R}^{d}\to{\mathscr{M}}_{d}(\mathbb{R})$, it is defined as
@@
This operator is called the (semiclassical) Kramers-Fokker-Planck (KFP) operator and takes the form
@@
One of the pioneer work studying $P_{KFP}$ is [19] in which the authors determined estimates regarding the return to equilibrium for the semigroup $e^{-tP_{KFP}}$ for certain potentials that are homogeneous near infinity. This result has then been generalized for example in [27, 2]. In [20] the authors derived a rough localization of the small eigenvalues of $P_{KFP}$ in the semiclassical limit $h\to 0$. More precise asymptotics have been obtained in [17, 18]. Note that these work are not restricted to the KFP operator, the authors considered a wider class of supersymmetric non-self-adjoint operators. Then, this supersymmetric assumption has been relaxed in [5]. The goal of this paper is to obtain a similar result for a KFP operator with general coefficients that does not fit the assumptions of [17, 18, 5].
@@ 1.2. Main results
However, the kinetic structure of $P$ is very important and yields powerful tools to palliate [17]. We can make use of the notion of hypocoercivity. This notion, was initiated in [21], and has been further developed in [33, 7] and references therein. Here we are in a setting which is very convenient to use a semiclassical version of [9] and we shall follow this strategy to obtain the desired resolvent estimates.
Throughout the paper, we shall consider a potential satisfying the following assumption.
@@
We consider the partition ${\mathcal{U}}={\mathcal{U}}^{odd}\sqcup{\mathcal{U}}^{even}$ where $x^{*}\in{\mathcal{U}}^{even}\iff\forall i,\ \nu_{i}^{x^{*}}\in 2\mathbb{N}$. Now we shall say that $x^{*}\in{\mathcal{U}}^{even}$ is of index $j\in\operatorname{[\negthinspace[}0,d\operatorname{]\negthinspace]}$ if $\sharp\{i\ |\ t_{i}^{x^{*}}<0\}=j$ and therefore ${\mathcal{U}}^{even}=\bigsqcup_{j=0}^{d}{\mathcal{U}}^{(j)}$ where ${\mathcal{U}}^{(j)}$ is the set of critical points of $V$ with even order in each direction and of index $j$, we also denote $n_{0}=\sharp{\mathcal{U}}^{(0)}$ the number of minima of $V$.
Notice that the critical points of $f$ are exactly the $(x^{*},0)$, with $x^{*}$ a critical point of $V$. Moreover, because $|\Sigma v|^{2}$ is quadratic and convex, we have the same partition ${\mathcal{U}}={\mathcal{U}}^{odd}\sqcup{\mathcal{U}}^{even}$ and the critical points have the same index when defined. In the following we will identify those two and use $x^{*}$ instead of $(x^{*},0)$ where it is clear which one we are really talking about ($x^{*}$ will mostly be denoted either $\mathbf{m}$ if of index $0$ or $\mathbf{s}$ if of index $1$).
@@
There exist $g_{1},g_{2}>0$ such that for all $x\in\mathbb{R}^{d}$,
@@
In the following we denote
(1.16) $\overline{\nu}=\max_{i,x^{*}}\nu_{i}^{x^{*}}.$
It allows us to introduce the function that will control the spectral gap of $P$,
@@
Suppose Assumptions 1, 2, 5, 6 and 9 hold true. There exist $h_{0}>0$, $c_{0},c_{1},c>0$, such that for all $h\in]0,h_{0}]$, there exists $G_{h}$ subspace of $L^{2}(\mathbb{R}^{d+d^{\prime}})$ of dimension $n_{0}$ such that
@@
then there exist $\lambda_{\mathbf{m}}(h)\in\mathbb{C}$ for all $\mathbf{m}\in{\mathcal{U}}^{(0)}$ such that $\sigma(P)\cap\{|\operatorname{Re}z|\leq c_{0}g(h)\}=\{\lambda_{\mathbf{m}}(h),\mathbf{m}\in{\mathcal{U}}^{(0)}\}$ counted with multiplicity, and for all $\mathbf{m}\in{\mathcal{U}}^{(0)},\ |\lambda_{\mathbf{m}}(h)|\leq e^{-c/h}$. Finally (still under (1.18)), for all $0<c_{0}^{\prime}<c_{1}$,
@@
For the rough localization of Theorem 1, we only need vague constructions around the critical points of $V$ in order to have Proposition 2.1. In order to obtain sharp asymptotics, we have to refine our constructions. It requires the introduction of the topological definitions we recall in Subsection A.2.
@@
This assumption comes from [6, (1.7)] and [11, Assumption 3.8]. Then it changed to become [18, Hypothesis 5.1] when finally reaching the form of [24, Assumption 4].
@@
The function $g$ defined in (1.17) is no worse than polynomial. That is, there exist $h_{0},c>0$ such that for all $h\in(0,h_{0}]$,
@@
Suppose the hypotheses of Theorem 1, Assumptions 7, 8, and (Gener) hold true. Suppose moreover that Proposition 4.1 apply. There exist $h_{0},\gamma>0$ such that for all $h\in\ ]0,h_{0}]$, one has $\lambda(\underline{\mathbf{m}},h)=0$ and for all $\mathbf{m}\neq\underline{\mathbf{m}}$, $\lambda(\mathbf{m},h)$ satisfies the following Eyring–Kramers type formula
@@
As mentioned in the paragraph after (Gener), it seems that this assumption is not required in order to obtain sharp results although we did no computation without it. For Assumption 8, we observe that it is very commonly satisfied and all examples in this article satisfy it.
@@ 1.3. Metastability
Furthermore, we can describe the metastable behavior of the solutions of (1.19).
@@
In the setting of Theorem 2, let $S_{1}\leq\cdots\leq S_{p+1}=+\infty$ denote the non-decreasing sequence of the $S(\mathbf{m})$’s defined in (A.1) such that if $S_{k}=S_{k+1}$, then $\mu_{k}<\mu_{k+1}$ and let $\Pi^{\leq}_{k}$ be the spectral projector of $P$ associated with its eigenvalues of modulus of order less than $h^{\mu_{k}}e^{-2S_{k}/h}$. For two positive functions $t_{\pm}(h)$ such that $t_{-}(h)=O(h^{\infty})$ and $t_{+}^{-1}(h)=o(h^{c-1}|\ln h|^{-1})$, we define the times
$t_{0}^{+}=t_{+}(h)\ \text{ and }\ \forall 1\leq k\leq p+1,\ t^{\pm}_{k}=t_{\pm}(h)h^{1-\mu_{k}}e^{2S_{k}/h}$
@@
We can link this with the stochastic process solving (1.5). Under mild hypotheses, one can show that the probability density $\rho(t,\cdot)$ of the process $(X_{t})_{t}$ solution of (1.5) is solution to the problem
@@ Acknowledgements
Then, we want to obtain precise Eyring-Kramers laws for our operator. We did not manage to obtain such results in broad generality, but we could prove these in several prescribed situations. Section 4 focuses on the local constructions of the WKB method in the spirit of [5], while the purpose of Section 5 is to glue these local constructions to obtain globally defined cutoffs. This leads to the proof of Theorem 2.
@@ 2. Hypocoercive estimates
The following assumption will help us prove some bounds on $A$ for the hypocoercivity result.
@@
    $\forall(x,v)\in\mathbb{R}^{d+d^{\prime}}\ \int\alpha(x,\sqrt{h}v)e^{-2|\Sigma v|^{2}}\ dv=0$.
@@
    where we denote $J_{x}\alpha$ the Jacobian matrix of $\alpha$ with respect to the variable $x$ (and likewise for $J_{v}\alpha$) and $\Delta_{v}\alpha$ the vector $(\Delta_{v}\alpha_{i})_{i}$.
@@ 2.1. Proof of Theorem 1
With the expression of $\delta(h)$ and Lemma 2.4, let us recall (1.17), for $h>0$, $g(h)$ is defined as
$g(h)=\frac{h}{1+h^{\frac{4}{\overline{\nu}}-2}\Big(\frac{g_{2}(h)}{g_{1}(h)}\Big)^{3}+h^{\frac{1}{\overline{\nu}}}g_{1}(h)^{-\frac{1}{2}}}.$
@@ 5.1. Proof of Theorem 2 and graded matrices
$e^{-tP/h}=\Pi_{0}+O(e^{-t\underset{\mathbf{m}\neq\underline{\mathbf{m}}}{\min}\operatorname{Re}(\lambda(\mathbf{m},h))(1-Ch)/h}),$
@@
which is a bounded operator, we have $\Pi_{0}=e^{-tP/h}\Pi_{0}$ and it suffices to show that
@@
and $D_{p+1}=D(0,R^{-1}h^{\mu_{p}}e^{-2S_{p}/h})$. For $R$ fixed large enough and every $h$ small enough, each exponentially small eigenvalue of $P$ belongs to exactly one of the disjoint sets $D_{k}$ from Theorem 2. Moreover, $\partial D_{k}$ is at distance of order $h^{\mu_{k}}e^{-2S_{k}/h}$ (resp. $h^{\mu_{p}}e^{-2S_{p}/h}$) from the spectrum of $P$ for $k\in\operatorname{[\negthinspace[}1,p\operatorname{]\negthinspace]}$ (resp. $k=p+1$). Using the resolvent estimate of $P$ on the image of $\Pi_{\mathscr{C}}$ given by Theorem 3 and (5.18) to control the contribution on the image of $1-\Pi_{\mathscr{C}}$, we get
@@
using that $\mu_{k}\geq 1$. On the contrary, for $t_{k}^{+}\leq t$, (5.21) and $e^{-t_{k}^{+}h^{\mu_{k}-1}e^{-2S_{k}/h}/R}=O(h^{\infty})$ give
@@
Lastly, $e^{-tP/h}\Pi_{p+1}=\Pi_{p+1}$ since $\Pi_{p+1}$ is the rank-one spectral projector associated with the eigenvalue $0$. On the other hand, since $e^{-\varepsilon h^{c-1}t_{0}^{+}}=O(h^{\infty})$, (5.20) becomes